OpenStudy (anonymous):

A customer went to a garden shop and bought some potting soil for 12.50 and 6 shrubs. If the total bill was $71.00, how much did each shrub cost if they all cost the same? Write and solve an equation Show your work. HELP PLEASE MEDAL WILL BE GIVENN

OpenStudy (dtan5457):

12.50+6x=71 6x=71-12.50 6x=58.5 9.75 Each shrub cost 9.75

OpenStudy (dtan5457):

Got it?

OpenStudy (anonymous):

Yes thank you.